Grandpa Gump is 63 years old.His age is 2 years less than 5 times the age of Billy Gump.How old is Billy?
Answer by TakeATuition.com(57)

Let The age of billy be x
Grandpa Gump =63 years
grandpa gump is 2 years less then 5 times age of billy =5x-2
therefore 5x-2=63 ......(1)equation
5x-2+2=63+2 adding +2 on both sides
5x=63+2
5x=65
5x/5=65/5 divding both sides by 5
x=13
there fore age of billy is 13 years we can cross check it by substituting its value in equation(1)
5x-2=63
5(13)-2=63
65-2=63
63=63 as value on left hand side is equal ot the value on right hand side therefore our answer is correct